# Optimal design of unimorph-type cantilevered piezoelectric energy harvesters using level set-based topology optimization by considering manufacturability

### v1

@@ 5.2 Optimal configuration subjected to a minimum output voltage constraint
Herein, the efficacy of the introduced constraint with respect to output voltage is critically examined. Consistently across all scenarios, the substrate-dependent constraint was enforced utilizing the fictitious physical function $p$. Moreover, the material domain was characterized by dual-level set functions, $\phi_{pe}$ and $\phi_{sb}$, with the parameter $\tau_{z}$ assigned a value of $1.0\times 10^{-2}$. Table 3 enumerates the parameters for the minimum output voltage constraint under various conditions, alongside their respective objective functions $F_{k}$, $F_{\omega}$, and output voltage $V_{E}$. The optimal configurations for each scenario are graphically represented in Figure 5.
@@
As shown in Table 3, the output voltage $V_{E}$ obtained under condition (j) exceeded the stipulated minimum voltage constraint. However, under the other conditions, the output voltages were aligned exactly with their respective constraints highlighted the effectiveness and precision of the minimum output voltage constraint. Furthermore, as the output voltage constraint values increased, there was a concomitant reduction in the volume of the piezoelectric material domain, represented by $\Omega_{pe}$. As elucidated by Eq. (12), this reduction in the volume of the piezoelectric material domain results in a diminished capacitance, which, in turn, amplifies the production voltage. Concomitantly, an increase in the $\bar{V}_{minV}$ value invariably leads to a corresponding increase in both objective functions $F_{k}$ and $F_{\omega}$. It is cogent to infer that imposing a minimum output voltage constraint invariably curtails the potential for optimizing the objective functions $F_{k}$ and $F_{\omega}$.
As shown in Figure 5, under conditions (k)–(q) with an active minimum output voltage constraint, the piezoelectric material domain was predominantly located near the fixed boundary and at the midpoint of the beam. Proximate to the fixed boundary of the beam, the aggregation of the piezoelectric material domain can be attributed to the stress concentration observed in the first-order eigenmode. The region encompassing the piezoelectric material domain located around the center of the beam was aligned with the zone of maximal displacement in the second-order eigenmodes.
@@ 6 Conclusion
In this study, we proposed a method for designing piezoelectric energy harvesters to improve their manufacturability. The proposed method derives both the substrate and piezoelectric designs that maximize the electromechanical coupling coefficient. This method allows the eigenfrequency of the device and minimum output voltage to be set to the desired values. Furthermore, the proposed method yields a design that can be manufactured using a microfabrication process. The results of this study are summarized as follows:
@@
    A topology optimization method was formulated to achieve the desired values for the eigenfrequency and minimum output voltage while concurrently maximizing the electromechanical coupling coefficient. An optimization algorithm was constructed for the numerical analysis.
@@
    Cross-sectional shape constraints and substrate-dependent constraints were formulated to improve manufacturability with microfabrication.
@@
    Several numerical examples were provided to demonstrate the utility and validity of the proposed method. The proposed method can provide solutions that satisfy the cross-sectional shape, substrate-depend, and minimum output voltage constraints.
@@ 6.1 Limitations and future work
Despite the promising results, this study has some limitations. As the constraints became more stringent, the value of the objective function deteriorated. This indicates that the proposed method may struggle to provide solutions with significant electrical-mechanical coupling coefficients under highly restrictive conditions.
@@ Appendix A Computation of the output voltage using the modal superposition method
In this study, output voltage was computed using the modal superposition method. In A.1, we explain the method used to solve the eigenvalue problem using the finite element method(FEM). In A.2, we explain the method used to yield the output voltage using the modal superposition method based on the results obtained in A.1.
@@ A.1 Eigenvalue and eigenvector determination
In this subsection, we elucidate an approach to solve the eigenvalue problem for piezoelectric materials using the FEM. Under open-circuit conditions, the relationships can be expressed in the form of finite element equations, as described below.
$(\bm{K}+j\omega\bm{R}-\omega_{oc}^{2}\bm{M})\bm{u}+\bm{P}\varphi$ $=0,$ (68)
@@
wherein $j$ denotes the imaginary unit; $\bm{u}$ and $\varphi$ correspond to the displacement and electric potential under open-circuit conditions, respectively; and $\bm{P}$ and $\bm{G}$ denote the piezoelectric and dielectric matrices, respectively. Notably, $\bm{R}$ signifies the damping coefficient matrix inherent to the system. In the absence of an external charge, denoted by $\bm{q}$, and an external force, symbolized by $\bm{f}$, the potential $\varphi$ can be eliminated. This leads to:
@@ A.2 Evaluation of displacement and output voltage using the modal superposition method
Because the eigenvectors were normalized, it was necessary to consider the magnitude of displacement under forced vibration conditions to evaluate the output voltage. To achieve this, we utilized the modal superposition method to determine the displacement. First, the equation of motion is expressed as,
$\bm{M}\ddot{\bm{u}}+\bm{C}\dot{\bm{u}}+\bm{Ku}=\bm{f}.$ (75)
For undamped free vibrations, we determined the eigenfrequency $\omega_{ocr}$ and the eigenvector $\bm{u}_{oci}$. Using the modal transformation matrix $\bm{\Phi}$, we performed a coordinate transformation such that $\bm{u}=\bm{\Phi}\bm{q}$. The modal transformation matrix $\bm{\Phi}$ is given by
$\bm{\Phi}=\left[\bm{u}_{oc1},\bm{u}_{oc2},\bm{u}_{oc3},\cdots,\bm{u}_{ocn}\right].$ (76)
Multiplying Eq. (75) on the left with $\bm{\Phi}^{T}$, we obtain:
$\bm{\Phi}^{T}\bm{M}\bm{\Phi}\ddot{\bm{q}}+\bm{\Phi}^{T}\bm{C}\bm{\Phi}\dot{\bm{q}}+\bm{\Phi}^{T}\bm{K}\bm{\Phi}\bm{q}=\bm{\Phi}^{T}\bm{f}.$ (77)
Through normalization, we transform each coefficient term as,
$\bm{\Phi}^{T}\bm{M}\bm{\Phi}$ $=\bm{I},$ (78)
$\bm{\Phi}^{T}\bm{C}\bm{\Phi}=\bm{D}$ $=\left[\begin{array}[]{llll}2\zeta_{1}\omega_{1}&0&\cdots&0\\ 0&2\zeta_{2}\omega_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&2\zeta_{n}\omega_{n}\end{array}\right],$ (83)
$\bm{\Phi}^{T}\bm{K}\bm{\Phi}$ $=\left[\begin{array}[]{llll}\omega_{1}^{2}&0&\cdots&0\\ 0&\omega_{2}^{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\omega_{n}^{2}\end{array}\right],$ (88)
$\bm{\Phi}^{T}\bm{f}$ $=\left[\begin{array}[]{llll}F_{1}(t)&0&\cdots&0\\ 0&F_{2}(t)&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&F_{n}(t)\end{array}\right].$ (93)
Assuming that the damping is small and the off-diagonal terms of the damping matrix are negligible compared with the diagonal terms, the transformation allowed us to express the motion Eq. (75) as $n$ independent equations for each eigenmode as,
@@
From the above , displacement is yielded as follows:
@@
Incorporating this $\bm{u}$ into Eq. (70) yields the potential $\varphi$. Notably, $\zeta_{i}=0.01$ for $i=1,2,\cdots n$. Finally, incorporating $\varphi$ into Eq. (10) yields the output voltage.
@@ Appendix B Sensitivity analysis
where the expressions within the brackets represent the directional derivatives of the functional. The optimality conditions, defined by Eqs. (97)– (112), revealed that the variables $\hat{\bm{u}}_{oc}$, $\hat{\bm{u}}_{sc}$, $\hat{\varphi}_{oc}$, $\hat{\omega}_{oc}$, and $\hat{\omega}_{sc}$ coincides with state the variables $\bm{u}_{oc}$, $\bm{u}_{sc}$, $\varphi_{oc}$, $\omega_{oc}$, and $\omega_{sc}$, respectively.
@@
From the above equations, the sensitivities for updating $\phi_{pe}$ and $\phi_{sb}$ can be obtained as follows:
$\begin{split}{F}^{\prime}_{pe}=\frac{\partial L_{pe}}{\partial\phi_{pe}}=&c_{ocpe}\left\{\frac{\partial a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})}{\partial\phi_{pe}}+\frac{\partial b_{oc}(\phi_{pe},\bm{u}_{oc},\varphi_{oc},\varphi_{oc})}{\partial\phi_{pe}}\right\}\\ &+c_{scpe}\left\{\frac{\partial a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})}{\partial\phi_{pe}}\right\}+\lambda,\\ =&c_{ocpe}\left\{a_{oc}(1,\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})+b_{oc}(1,\bm{u}_{oc},\varphi_{oc},\varphi_{oc})\right\}\\ &+c_{scpe}a_{sc}(1,\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})+\lambda,\end{split}$ (119)
$\begin{split}{F}^{\prime}_{sb}=\frac{\partial L_{sb}}{\partial\phi_{sb}}=&c_{ocsb}\left\{\frac{\partial a_{oc}(\phi_{pe},\phi_{sb},\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})}{\partial\phi_{sb}}+\frac{\partial b_{oc}(\phi_{pe},\bm{u}_{oc},\varphi_{oc},\varphi_{oc})}{\partial\phi_{sb}}\right\}\\ &+c_{scsb}\left\{\frac{\partial a_{sc}(\phi_{pe},\phi_{sb},\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})}{\partial\phi_{sb}}\right\},\\ =&c_{ocsb}\left\{a_{oc}(\phi_{pe},1,\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})\right\}\\ &+c_{scsb}\left\{a_{sc}(\phi_{pe},1,\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})\right\}.\end{split}$ (120)

### v2

@@ 5.2 Optimal configuration subjected to a minimum output voltage constraint
Herein, the efficacy of the introduced constraint with respect to output voltage is critically examined. Consistently across all scenarios, the substrate-dependent constraint is enforced utilizing the fictitious physical function ${\xi}$. Moreover, the material domain is characterized by two level set functions, ${\phi_{p\psi}}$ and ${\phi_{s\psi}}$, with the parameter $\tau_{z}$ assigned a value of $1.0\times 10^{-2}$. Table 3 enumerates the parameters for the minimum output voltage constraint under various conditions, alongside their respective objective functions $F_{k}$, $F_{\omega}$, and output voltage $V_{E}$. The optimal configurations for each scenario are graphically represented in Figure 11.
@@
As shown in Table 3, the output voltage $V_{E}$ obtained under condition (j) exceeds the stipulated minimum voltage constraint. However, under the other conditions, the output voltages are aligned exactly with their respective constraints, highlighting the effectiveness and precision of the minimum output voltage constraint. Furthermore, as the output voltage constraint values increase, there is a concomitant reduction in the volume of the piezoelectric material domain, represented by $\Omega_{pe}$. As elucidated by Eq. (26), this reduction in the volume of the piezoelectric material domain results in a diminished capacitance, which, in turn, amplifies the production voltage. Concomitantly, an increase in the $\bar{V}_{minV}$ value invariably leads to a corresponding increase in both objective functions $F_{k}$ and $F_{\omega}$. It is cogent to infer that imposing a minimum output voltage constraint invariably curtails the potential for optimizing the objective functions $F_{k}$ and $F_{\omega}$.
As shown in Figure 11, under conditions (k)–(q) with an active minimum output voltage constraint, the piezoelectric material domain is predominantly located near the fixed boundary and at the midpoint of the beam. Close to the fixed boundary of the beam, the aggregation of the piezoelectric material domain can be attributed to the stress concentration observed in the first-order eigenmode. The region encompassing the piezoelectric material domain located around the center of the beam is aligned with the zone of maximal displacement in the second-order eigenmodes.
@@ 6 Conclusion
In this study, we propose a method for designing piezoelectric energy harvesters while improving their manufacturability. The proposed method derives both the substrate and piezoelectric designs that maximize the electromechanical coupling coefficient. This method allows the eigenfrequency of the device and minimum output voltage to be set to the desired values. Furthermore, the proposed method yields a design that can be manufactured using a microfabrication process. The results of this study are summarized as follows:
@@
    A topology optimization method is formulated to achieve the desired values for the eigenfrequency and minimum output voltage while concurrently maximizing the electromechanical coupling coefficient. An optimization algorithm is constructed for the numerical analysis.
@@
    Cross-sectional shape constraints and substrate-dependent constraints are formulated to improve manufacturability with microfabrication.
@@
    Several numerical examples are provided to demonstrate the utility and validity of the proposed method. The proposed method can provide solutions that satisfy the cross-sectional shape, substrate-dependent, and minimum output voltage constraints.
@@ 6.1 Limitations and future work
Despite the promising results, this study has some limitations. As the constraints become more stringent, the value of the objective function deteriorates. This indicates that the proposed method may struggle to provide solutions with significant electrical-mechanical coupling coefficients under highly restrictive conditions.
@@ Appendix A Computation of the output voltage using the modal superposition method
In this study, the output voltage is computed using the modal superposition method. In A.1, we explain the method used to solve the eigenvalue problem using the finite element method (FEM). In A.2, we explain the method used to yield the output voltage using the modal superposition method based on the results obtained in A.1.
@@ A.1 Eigenvalue and eigenvector determination
In this subsection, we elucidate an approach to solve the eigenvalue problem for piezoelectric materials using the FEM. Under open-circuit conditions, the relationships can be expressed in the form of finite element equations, as described below [4, 24].
$(\bm{K}-\omega_{oc}^{2}\bm{M})\bm{u}+\bm{P}\varphi$ $=0,$ (112)
@@
wherein $\bm{u}$ and $\varphi$ correspond to the displacement and electric potential under open-circuit conditions, respectively , and $\bm{P}$ and $\bm{G}$ denote the piezoelectric and dielectric matrices, respectively. In the absence of an external charge, denoted by $\bm{q}$, and an external force, symbolized by $\bm{f}$, the potential $\varphi$ can be eliminated. This leads to:
@@ A.2 Evaluation of displacement and output voltage using the modal superposition method
In this subsection, we describe the process of obtaining the displacement field using the modal superposition method, utilizing the eigenvectors obtained in the previous subsection [4]. To evaluate the output voltage, it is necessary to determine the displacement field. Since the eigenvectors are normalized, we need to consider the magnitude of the response for each mode to accurately determine the overall displacement. Therefore, we employ the modal superposition method to determine the displacement field necessary for this evaluation.
To obtain the displacement field under forced vibration conditions, we initially express the equation of motion as:
$\bm{M}\ddot{\bm{u}}+\bm{C}\dot{\bm{u}}+\bm{K}\bm{u}=\bm{f}.$ (119)
In the previous subsection, we discussed the general eigenvalue problem and denoted the solutions as $\bm{u}_{oc}$, $\omega_{oc}$, $\bm{u}_{sc}$, and $\omega_{sc}$. Using the modal transformation matrix $\bm{{\Lambda}}$, we perform a coordinate transformation such that $\bm{u}=\bm{{\Lambda}}\bm{q}$. The modal transformation matrix $\bm{{\Lambda}}$ and the modal amplitude vector $\bm{q}$ are given by:
$\bm{{\Lambda}}=$ $\left[\bm{u}_{oc1},\bm{u}_{oc2},\bm{u}_{oc3},\cdots,\bm{u}_{ocn}\right],$ (120)
${\bm{q}=}$ ${\left[{q}_{1},{q}_{2},{q}_{3},\cdots,{q}_{n}\right].}$ (121)
Multiplying Eq. (119) on the left with $\bm{{\Lambda}}^{T}$, we obtain:
$\bm{{\Lambda}}^{T}\bm{M}\bm{{\Lambda}}\ddot{\bm{q}}+\bm{{\Lambda}}^{T}\bm{C}\bm{{\Lambda}}\dot{\bm{q}}+\bm{{\Lambda}}^{T}\bm{K}\bm{{\Lambda}}\bm{q}=\bm{{\Lambda}}^{T}\bm{f}.$ (122)
Through normalization, we transform each coefficient term as follows:
$\bm{{\Lambda}}^{T}\bm{M}\bm{{\Lambda}}$ $=\bm{I},$ (123)
$\bm{{\Lambda}}^{T}\bm{C}\bm{{\Lambda}}=\bm{D}$ $=\left[\begin{array}[]{llll}2\zeta_{1}\omega_{1}&0&\cdots&0\\ 0&2\zeta_{2}\omega_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&2\zeta_{n}\omega_{n}\end{array}\right],$ (128)
$\bm{{\Lambda}}^{T}\bm{K}\bm{{\Lambda}}$ $=\left[\begin{array}[]{llll}\omega_{1}^{2}&0&\cdots&0\\ 0&\omega_{2}^{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\omega_{n}^{2}\end{array}\right],$ (133)
$\bm{{\Lambda}}^{T}\bm{f}$ $=\left[\begin{array}[]{llll}F_{1}(t)&0&\cdots&0\\ 0&F_{2}(t)&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&F_{n}(t)\end{array}\right].$ (138)
where $\zeta_{i}$ is the modal damping ratio for the $i$th mode. In this transformation, we assume that the damping is small and the off-diagonal terms of the damping matrix are negligible compared with the diagonal terms. This transformation allows us to express the motion Eq. (119) as $n$ independent equations , one for each eigenmode as follows:
@@
From the above, the displacement is obtained as:
@@
Incorporating this $\bm{u}$ into Eq. (114) yields the potential $\varphi_{n}$. Notably, $\zeta_{i}=0.01$ for $i=1,2,\cdots n$. Finally, incorporating $\varphi_{n}$ into Eq. (24) yields the output voltage.
@@ Appendix B Sensitivity analysis
where the expressions within the brackets represent the directional derivatives of the functional. The variables denoted with a hat, such as $\hat{\bm{u}}$, represent the variables with respect to which the variational analysis is performed and should not be confused with the state variables, which are the solutions of the state equations. This distinction is made to clearly differentiate between the variables used in the sensitivity analysis and the state variables.
@@
From the above equations, the sensitivities for updating ${\phi_{p\psi}}$ and ${\phi_{s\psi}}$ can be obtained as follows:
$\begin{split}{F}^{\prime}_{pe}=\frac{\partial L_{pe}}{\partial{\phi_{p\psi}}}=&c_{ocpe}\left\{\frac{\partial a_{oc}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})}{\partial{\phi_{p\psi}}}+\frac{\partial b_{oc}({\phi_{p\psi}},\bm{u}_{oc},\varphi_{oc},\varphi_{oc})}{\partial{\phi_{p\psi}}}\right\}\\ &+c_{scpe}\left\{\frac{\partial a_{sc}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})}{\partial{\phi_{p\psi}}}\right\}+\lambda,\\ =&c_{ocpe}\left\{a_{oc}(1,{\phi_{s\psi}},\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})+b_{oc}(1,\bm{u}_{oc},\varphi_{oc},\varphi_{oc})\right\}\\ &+c_{scpe}a_{sc}(1,{\phi_{s\psi}},\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})+\lambda,\end{split}$ (164)
$\begin{split}{F}^{\prime}_{sb}=\frac{\partial L_{sb}}{\partial{\phi_{s\psi}}}=&c_{ocsb}\left\{\frac{\partial a_{oc}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})}{\partial{\phi_{s\psi}}}+\frac{\partial b_{oc}({\phi_{p\psi}},\bm{u}_{oc},\varphi_{oc},\varphi_{oc})}{\partial{\phi_{s\psi}}}\right\}\\ &+c_{scsb}\left\{\frac{\partial a_{sc}({\phi_{p\psi}},{\phi_{s\psi}},\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})}{\partial{\phi_{s\psi}}}\right\},\\ =&c_{ocsb}\left\{a_{oc}({\phi_{p\psi}},1,\omega_{oc},\bm{u}_{oc},\bm{u}_{oc},\varphi_{oc})\right\}\\ &+c_{scsb}\left\{a_{sc}({\phi_{p\psi}},1,\omega_{sc},\bm{u}_{sc},\bm{u}_{sc})\right\}.\end{split}$ (165)